a curve y=1 - 4/(2x+1)square. the curve intersect the x- axes at A . the normal to the curve at A intersects the y- axes at B.
1) obtain expressions for dy/dx and( upper and lower limit ) |ydx
2) find showing all necessary working find all the area under the curve and normal (full)
1
Expert's answer
2020-04-27T18:28:42-0400
"y=1-{4 \over (2x+1)^2}"
"y=0=>1-{4 \over (2x+1)^2}=0=> (2x+1)^2=4"

"2x+1=-2\ or \ 2x+1=2"

"x=-1.5\ or\ x=0.5"

Point "A(0.5,0)"


"y'=(1-{4 \over (2x+1)^2})'=-{4(-2) \over (2x+1)^3}(2x+1)'={16 \over (2x+1)^3}"

"y'0.5)={16 \over (2(0.5)+1)^3}=2"

Fnd the equation of the normal at the point "A(0.5,0)"


"slope=m=-{1 \over y'(0.5)}=-{1 \over 2}=-0.5"

"0=-0.5(0.5)+b=>b=0.25"

The equation of the normal is "y=-0.5x+0.25."

Point "B(0,0.25)"


"S=\displaystyle\int_{0}^{0.5}\big(-0.5x+0.25-(1-{4 \over (2x+1)^2})\big)dx="

"=\big[-0.25x^2-0.75x-{2 \over 2x+1}\big]\begin{matrix}
 0.5 \\
 0
\end{matrix}="

"=-0.0625-0.375-1+0+0+2=0.5625(units^2)"

The area is "0.5625" square units.